The following data are taken from the financial statements of Sigmon Inc. Terms of all sales are 2/10, n/45 20Y3 20Y2 20Y1 Accounts receivable, end of year Sales on account 725,000 650,000 4,687,500 $600,000 5,637,500

Horizontally analyze the financial data in Exercise 17-9. (b) Calculate how many days it took Sigmon Inc. to collect its recieivables (i.e. Sigmon's average collection period) in 20Y3. Show your work.
Answer #1

Receivable turnover ratio =Sales revenue/Average account receivable

Average account receivable=(725000+650000)/2=$687500

Receivable turnover ratio=5637500/687500=8.2 time

Average collection Period =365/Receivable turnover ratio=365/8.2=44.51 Days